OpenStudy (anonymous):

hi pls. help me with this. any number divided by ∞=0 is it correct? then what about any number added and subtracted to ∞?

OpenStudy (anonymous):

it will be ∞ ∞ + x = ∞ ∞ - x = ∞

OpenStudy (anonymous):

x/ ∞ = 0 and x/0 = ∞

OpenStudy (anonymous):

wow! tnx

OpenStudy (anonymous):

∞ is not a number. You can't multiply or divide by it, nor can add or subtract to anything. All the problems stated above should be evaluated by limits.

OpenStudy (anonymous):

well then how should i solve limit of 1/1+2E1/x as x→-∞?

OpenStudy (anonymous):

Sorry the first limit doesn't actually exist, let me fix it.

OpenStudy (anonymous):

It should be noted that there is an extended version of the real number field with positive and negative infinity. Yet they are still not considered numbers. Instead of what noru said you could express the intuition with these two limits: Let c be some constant in the real number field \[\lim_{x \rightarrow 0}\left|\frac{c}{x}\right| = \infty\] \[\lim_{x \rightarrow \infty} \frac{c}{x} = 0\]

OpenStudy (anonymous):

In this case it doesnt matter what the sign is as x goes to infinity because it will be 0 either way.

OpenStudy (anonymous):

=1/1+2to the power of 1/-∞ =1/1+∞=1/∞ =0 is it correct?

OpenStudy (anonymous):

Is this what you mean? \[\lim_{x \to -\infty}\frac{1}{1+2^{1/x}}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

but the limit of 1/x as x approaches infinity is 0 2^0 = 1 so its\[\frac{1}{1+1} = \frac{1}{2}\]

OpenStudy (anonymous):

Think about it this way, as x gets very large, the power 2 is raised to is very small.

OpenStudy (anonymous):

ahm... i did not understand?yes 1/x as x approaches infinity is o.then did it come up to 1 when it is 2 to the power of 0?

OpenStudy (anonymous):

Anything to the power of 0 is 1.

OpenStudy (anonymous):

oh yup tnx